Simon operates an HST-registered mobile glass repair service in Ontario. His service revenue for the year...

  1. Simon operates an HST-registered mobile glass repair service in Ontario. His service revenue for the year is $28 000. His HST-taxable purchases amounted to $4000. Simon does not use the Quick Method for claiming ITCs. By calculating the difference between the HST he collected and the HST he paid, determine Simon’s HST remittance to the CRA.
Answer #1
Province terriroty Ontario
Assuming amount is exclusive of HST taxes
Tax on sales
Sale amount 28000 Before tax
HST
Fed 5% 1400 =28000*0.05
Prov 8% 2240 =28000*0.08
3640
After tax amount 31640
Tax on purchases
Purchase price 4000 Before tax
Fed 5% 200 =4000*0.05
Prov 8% 320 =4000*0.08
520
After tax amount 31640
HST Payable 3640
HST Paid (ITC) 520
HST remittable 3120
Assuming amount is inclusive of HST taxes
Tax on sales
Sales Price 28000 After tax
Fed 5% 1238.94 =28000*5%/113%
Prov 8% 1982.3 =28000*8%/113%
3221.24
Before tax amount 24778.76
Tax on purchases
Purchase price 4000 After tax
Fed 5% 176.99 =4000*5%/113%
Prov 8% 283.19 =4000*8%/113%
460.18
Before tax amount 3539.82
HST Payable 3221.76
HST Paid (ITC) 460.18
HST remittable 2761.58